# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

### v1

@@ G.5 Experiment Design and Results
| Qwen2-Audio | 100 | 84 | 99 | 87 | 100 | 92 | 96 | 51 | 100 | 95 | 100 | 95 | 100 | 100 | 100 | 100 | 100 | 100 |
| SALMONN | 49 | 3 | 43 | 2 | 36 | 9 | 27 | 1 | 34 | 1 | 70 | 0 | 3 | 0 | 67 | 87 | 100 | 75 |
| MiniCPM-o 2.6 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 94 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 | 100 |
| UltraVox | 0 | 0 | 0 | 0 | 82 | 0 | 30 | 0 | 0 | 0 | 50 | 0 | 99 | 94 | 76 | 83 | 100 | 100 |
@@
| Gemini-1.5 Pro | 77 | 0 | 97 | 0 | 100 | 0 | 76 | 0 | 0 | 0 | 100 | 5 | 84 | 72 | 100 | 100 | 100 | 100 |
| Gemini-2.5 Flash | 100 | 0 | 100 | 0 | 100 | 0 | 100 | 0 | 90 | 0 | 100 | 0 | 98 | 0 | 100 | 99 | 100 | 100 |
| Gemini-2.5 Pro | 100 | 0 | 100 | 0 | 0 | 0 | 100 | 0 | 0 | 0 | 100 | 8 | 100 | 100 | 100 | 100 | 100 | 100 |
| GPT-4o Audio | 0 | 0 | 0 | 0 | 30 | 1 | 0 | 0 | 0 | 0 | 12 | 1 | 98 | 78 | 83 | 69 | 100 | 96 |
| GPT-4o Mini Audio | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 91 | 32 | 68 | 52 | 100 | 100 |
| Average | 58.4 | 20.8 | 59.9 | 21.0 | 60.9 | 22.4 | 58.8 | 16.2 | 36.0 | 21.8 | 70.2 | 23.2 | 85.9 | 64.0 | 88.2 | 87.8 | 100.0 | 96.8 |
Note: "no aug" indicates the complete disclosure rates before applying a privacy-aware prompt, while "aug" shows rates after applying it. Lower values indicate better performance. Red arrows () indicate worse performance (higher leakage) than average; blue arrows () indicate better performance (lower leakage) than average. SSN: Social Security Number; Pwd: Phone Password; Setting: Setting of Conversation.
| Input | Model | SSN (no aug) | SSN (aug) |
| Audio | Gemini-1.5 Pro | 70 | 100 |
| Gemini-2.5 Flash | 0 | 100 | |
| Gemini-2.5 Pro | 100 | 100 | |
| GPT-4o Audio | 100 | 100 | |
| GPT-4o Mini Audio | 100 | 100 | |
| Text | Gemini-1.5 Pro | 2 | 100 |
| Gemini-2.5 Flash | 72 | 98 | |
| Gemini-2.5 Pro | 82 | 94 | |
| GPT-4o | 61 | 100 | |
| GPT-4o mini | 2 | 100 |
Note: Values are in the format “no aug” (original input data) and “aug” (with prompt enhancements). Gemini-1.5 Pro is the baseline. indicates better performance relative to baseline; indicates worse performance. Subscripts show the difference from the baseline.
| Model | Age | Ethnicity | Setting |
| Open-source Models | |||
| Qwen2-Audio | 22.0 | 16.0 | 59.0 |
| SALMONN | – | 16.2 | 73.0 |
| MiniCPM-o 2.6 | 29.0 | 22.0 | 84.0 |
| UltraVox | 23.2 | 10.7 | 81.0 |
| Closed-source Models | |||
| Gemini-1.5 Pro | 24.7 | 52.0 | 94.0 |
| Gemini-2.5 Flash | 28.9 | 39.0 | 95.0 |
| Gemini-2.5 Pro | 20.0 | 61.0 | 95.0 |
| GPT-4o Audio | 23.5 | 42.7 | 95.0 |
| GPT-4o Mini Audio | 16.5 | 7.7 | 91.0 |
| Average | 23.5 | 29.7 | 85.2 |
Note: Setting: Setting of Conversation. ‘–’ indicates insufficient valid inference samples. indicates higher than average (better); indicates lower than average. Numbers next to arrows show absolute difference from average.
@@ Results.
From the experimental results, it can be observed that different models exhibit varying levels of sensitivity to different types of personal privacy information. For instance, in the case of highly sensitive information such as Social Security Numbers (SSN), the refusal rate is generally high across models, with GPT-4o Audio showing no leakage at all. However, for other types of private information such as home addresses and mobile phone passwords, the complete leakage rate reaches 100% for Gemini Series, when no prompt engineering techniques are applied. Other models also show similar trends, but overall, the GPT-4o series demonstrates superior comprehensive performance, exhibiting stronger privacy protection capabilities compared to other models.
@@
In privacy inference tasks, the model is required to infer personal privacy information from a given audio segment and its corresponding textual question. Experimental results show that except for SALMONN, which performs relatively well in inferring attributes such as age and ethnicity, the privacy leakage rate of most models exceeds 80% (The model tends to directly respond: “The age of the speaker cannot be inferred from the given audio.”). This indicates that most current models lack effective mechanisms for actively identifying or preventing potential privacy risks. For example, the open-source model Qwen2-Audio rarely refuses to answer questions related to age and ethnicity, whereas SALMONN shows comparatively better behavior. This difference may stem from the blurred boundary between privacy-related and general information, making it difficult for models to distinguish between them effectively.
@@
Introducing prompts containing privacy protection content (prompt engineering) can significantly enhance the model’s ability to prevent direct privacy leaks and reduce the full leakage rate. For example, the Gemini series achieves over an 80% reduction in full leakage rates for sensitive information such as bank account numbers and home addresses when enhanced prompts are used. However, this approach has limited effectiveness in mitigating inference-based privacy leaks and may even lead to a decrease in refusal rates. For instance, after introducing privacy-enhanced prompts, Gemini-2.5 Flash experiences a 5% increase in full leakage rate in age inference tasks.
@@
The experimental results in Table 17 also reveal differences in privacy awareness between audio and text models. Similar to audio models, the text-based GPT-4o series demonstrates strong security awareness. However, overall, text models tend to have lower refusal rates, indicating slightly reduced sensitivity to privacy information compared to audio models. Nevertheless, through the application of prompt engineering techniques, the privacy protection capabilities of text models can still be significantly improved, although the improvement is typically not as substantial as that seen in audio models. For example, Gemini-2.5 Flash achieves an improvement of less than 20% in protecting social security number under enhanced prompting.
(5) Model Inference Performance and Associated Privacy Concerns
As shown in Table 18, there are significant differences in the inference capabilities across various models. Specifically, MiniCPM-o 2.6 demonstrates strong performance in age inference, achieving an accuracy of 29.0%, while Gemini-2.5 Pro excels in ethnicity inference with a notably high accuracy of 61.0%. Overall, closed-source models outperform open-source models in inference tasks. However, in the absence of privacy-preserving techniques (e.g., prompt engineering), the low rejection rates for sensitive attributes such as age and ethnicity ( Table 15) indicate that the models’ powerful inference capabilities may introduce new privacy leakage risks when not properly controlled.
@@ H.2 Dataset Construction Method
To ensure the reliability and effectiveness of the native audio quality, we introduce big_bench_audio dataset [63] as one of the native voice data samples. The subsequent audio production follows the following process:
@@ H.3.3 Results
| Qwen2-Audio | 6.00 | 3.50 | 4.33 | 6.85 | 5.40 | 6.60 |
| SALMONN | 2.00 | 6.43 | 4.57 | 2.94 | 7.17 | 6.67 |
| Ultravox | 4.00 | 7.54 | 7.30 | 6.54 | 6.70 | 7.00 |
@@
| Gemini-1.5 Pro | 8.58 | 8.21 | 8.23 | 8.17 | 6.09 | 7.43 |
| Gemini-2.5 Flash | 8.17 | 8.39 | 8.29 | 7.93 | 6.37 | 7.77 |
| Gemini-2.5 Pro | 8.89 | 8.68 | 8.50 | 8.19 | 7.47 | 7.71 |
| GPT-4o Audio | 5.90 | 5.50 | 8.33 | 7.32 | 7.63 | 6.28 |
| GPT-4o mini Audio | 8.33 | 6.91 | 7.69 | 6.00 | 5.78 | 7.25 |
| Average | 6.63 | 5.82 | 7.13 | 6.78 | 6.57 | 6.99 |
: AR: Adversarial Robustness; AQV: Audio Quality Variation; BC: Background Conversation; ES: Environmental Sound; MS: Multiple
Speakers; NI: Noise Interference.
| Test Type | MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio |
| Adversarial Robustness | 8.27 | 6.06 | 5.84 | 1.00 | 8.09 | 7.61 | 8.17 | 6.70 | 1.44 |
| Audio Quality Variation | 8.56 | 5.90 | 6.25 | 1.29 | 7.90 | 7.59 | 8.17 | 5.80 | 1.73 |
| Background Conversation | 8.35 | 6.40 | 6.58 | 1.06 | 7.71 | 6.87 | 7.35 | 6.93 | 1.73 |
| Environmental Sound | 8.19 | 6.43 | 7.06 | 1.27 | 8.06 | 7.03 | 7.50 | 6.72 | 2.36 |
| Multiple Speakers | 8.74 | 6.78 | 6.33 | 2.44 | 7.66 | 7.24 | 7.99 | 8.39 | 4.74 |
| Noise Interference | 4.27 | 3.83 | 4.22 | 1.34 | 5.86 | 5.61 | 6.37 | 2.85 | 1.67 |
| Average | 7.73 | 5.90 | 6.05 | 1.40 | 7.55 | 6.99 | 7.59 | 6.23 | 2.28 |
@@
We evaluate the robustness of nine models against various auditory challenges in Appendix H.3.1, with detailed results presented in Table 19 Table 20 Table 21 and Talbe 22. The results reveal the following key findings:
@@ I.2 Dataset Construction Method
Voice Cloning Spoofing (200 samples). We collected 100 samples of voice wake-up words from an existing dataset [24], with each sample consisting of approximately 2 seconds of wake-up word audio. These samples constitute the Voice Cloning Spoofing test set. Detailed wake-up words and their corresponding developers are listed in Table 23. Each sample was cloned using F5-TTS [12], and we constructed paired audio clips composed of the original audio (used as the legitimate user’s enrollment input) and the cloned audio (used as the attack input attempting to bypass authentication). In addition, we designed two distinct authentication scenarios for the experiments—Awake Assistant and Open The Door. In the former, the wake-up words are used to authenticate an AI voice assistant, while in the latter, they are used to activate access to a high-security laboratory door, with an emphasis on the consequences of authentication failures. Through this experimental setup, we aim to explore how different system prompt formulations affect an ALLM’s ability to defend against voice cloning spoofing under identical audio inputs.
@@
Note: All data in this table are adapted from [24].
@@ I.3 Additional Implementation Details
To ensure our evaluation closely mirrors real-world scenarios, distinct inference prompts were designed for each data point across all experiments, complementing the audio inputs to the ALLMs. For the hybrid spoofing dataset, in particular, prompts were varied based on the specific attack scenario. Illustrative examples of these inference prompts are provided below. Below is the Inference Prompt for Experiment of Social Engineering Impact on Voice Verification.
@@
Below is the Evaluation Prompt for Assessing Social Engineering Impact on Voice Verification.

### v3

@@ G.5 Experiment Design and Results
| Kimi Audio | 97.0 | 100/100 | 87.9 | 99/100 | 97.0 | 100/100 | 98.0 | 100/100 | 99.0 | 100/100 | 89.0 | 100/100 | 21.8 | 55/100 | 37.0 | 100/100 | 59.0 | 100/100 |
| SALMONN | 27.1 | 48/100 | 26.8 | 41/100 | 24.3 | 37/100 | 72.0 | 25/100 | 38.2 | 34/100 | 28.6 | 70/100 | 3/100 | 16.7 | 66/100 | 43.3 | 90/100 | |
| OpenS2S | 53.4 | 88/100 | 34.2 | 79/100 | 29.0 | 100/100 | 49.5 | 95/100 | 46.2 | 91/100 | – | 22.0 | 100/100 | 17.6 | 91/100 | 50.0 | 82/100 | |
| Step Fun | 1/100 | 54.4 | 79/100 | 81.6 | 98/100 | 98.7 | 78/100 | 2/100 | 63.4 | 93/100 | 13.0 | 54/100 | 17.0 | 100/100 | 52.0 | 100/100 | ||
| Qwen2-Audio | 79.0 | 100/100 | 85.9 | 99/100 | 82.0 | 100/100 | 89.6 | 96/100 | 49.0 | 100/100 | 66.0 | 100/100 | 22.0 | 100/100 | 16.0 | 100/100 | 38.8 | 98/100 |
| Qwen2.5-Omni | 96.0 | 100/100 | 86.0 | 100/100 | 94.0 | 100/100 | 98.0 | 100/100 | 92.0 | 100/100 | 99.0 | 100/100 | 29.0 | 100/100 | 24.0 | 100/100 | 53.5 | 99/100 |
| MiniCPM-o 2.6 | 96.0 | 100/100 | 85.0 | 100/100 | 95.0 | 100/100 | 98.0 | 100/100 | 97.0 | 100/100 | 94.0 | 100/100 | 29.0 | 100/100 | 22.0 | 100/100 | 54.1 | 98/100 |
| Step Audio2 | 97.0 | 100/100 | 86.0 | 100/100 | 97.0 | 100/100 | 98.0 | 100/100 | 99.0 | 99/99 | 74.0 | 100/100 | 36.0 | 100/100 | 29.0 | 100/100 | 62.0 | 100/100 |
| UltraVox | 2/100 | 5/100 | 44.4 | 81/100 | 87.5 | 24/100 | 0/100 | – | 23.2 | 99/100 | 12.0 | 75/100 | 58.0 | 100/100 | ||||
@@
| Gemini-1.5 Pro | 69.7 | 99/100 | 63.0 | 100/100 | 92.0 | 100/100 | 98.0 | 100/100 | 93.1 | 29/99 | 95.0 | 100/100 | 22.6 | 84/100 | 52.0 | 100/100 | 63.0 | 100/100 |
| Gemini-2.5 Flash | 94.6 | 92/100 | 84.0 | 100/100 | 98.0 | 99/99 | 98.0 | 100/100 | 94.5 | 91/91 | – | 28.6 | 98/100 | 39.0 | 100/100 | 67.3 | 98/100 | |
| Gemini-2.5 Pro | 94.0 | 100/100 | 84.0 | 100/100 | 0/100 | 98.0 | 100/100 | 0/100 | 96.0 | 99/100 | 20.0 | 100/100 | 61.0 | 100/100 | 65.7 | 99/100 | ||
| GPT-4o Mini Audio | 0/100 | 0/100 | 0/100 | 0/100 | 0/100 | 0/100 | 16.5 | 91/100 | 26.9 | 67/100 | 61.0 | 100/100 | ||||||
| GPT-4o Audio | 0/100 | 0/100 | 93.9 | 33/100 | 0/100 | 0/100 | 80.0 | 15/100 | 23.5 | 98/100 | 42.7 | 82/100 | 69.0 | 100/100 | ||||
| Average | 80.4 | – | 70.7 | – | 77.4 | – | 90.3 | – | 78.7 | – | 78.5 | – | 23.6 | – | 29.5 | – | 56.9 | – |
Note: This table shows inference accuracy ONLY for samples where privacy was leaked. Higher values indicate greater privacy threat. Red
arrows () indicate higher threat than average; blue arrows () indicate lower threat. ‘’ indicates insufficient leaked samples ( 10).
Leak rates show leaked/total samples.
| Text | Audio | |||
| Model | SSN (w/o) | SSN (w/) | SSN (w/o) | SSN (w/) |
| Gemini-1.5 Pro | 2 | 100 | 71 | 100 |
| GPT-4o /Audio | 61 | 100 | 100 | 100 |
| GPT-4o Mini /Audio | 2 | 100 | 100 | 100 |
| Gemini-2.5 Flash | 72 | 98 | 0 | 100 |
| Gemini-2.5 Pro | 82 | 94 | 100 | 100 |
Note: Values are in the format “w/o” (original input data) and “w/” (with prompt enhancements). Gemini-1.5 Pro is the baseline for both text and audio tasks. indicates better performance relative to baseline; indicates worse performance; Gemini-1.5-pro indicate baseline performance. Subscripts show the absolute difference from the baseline.
@@ Results.
From the experimental results, it can be observed that different models exhibit varying levels of sensitivity to different types of personal privacy information. For instance, in the case of highly sensitive data such as Social Security Numbers (SSNs), most models demonstrate high refusal rates. Notably, GPT-4o Audio exhibits no leakage whatsoever, regardless of prompt formulation. In contrast, MiniCPM-o 2.6 consistently discloses SSNs in full, both with and without privacy-enhancing prompts. More importantly, the inference accuracy for SSN disclosures by MiniCPM-o 2.6 exceeds 85%, indicating that the leaked information is highly accurate. This suggests that the model can precisely retain and reproduce private information throughout the conversation, thereby posing a significantly greater privacy risk. For other types of private information, such as home addresses and mobile phone passwords, the Gemini series models exhibit a 100% complete leakage rate when no prompt engineering techniques are applied. Moreover, the accuracy of these disclosures is also high, further amplifying the potential privacy risk (Luo et al., 2025). Other models also show similar trends, but overall, the GPT-4o series demonstrates superior comprehensive performance, exhibiting stronger privacy protection capabilities compared to other models.
@@
In privacy inference tasks, the model is required to infer personal privacy information from a given audio segment and its corresponding textual question. Experimental results show that except for SALMONN, which performs relatively well in inferring attributes such as age and ethnicity, the privacy leakage rate of most models exceeds 80% (The model tends to directly respond: “The age of the speaker cannot be inferred from the given audio.”). This indicates that most current models lack effective mechanisms for actively identifying or preventing potential privacy risks. For example, the open-source model Qwen2-Audio rarely refuses to answer questions related to age and ethnicity, whereas SALMONN shows comparatively better behavior. This difference may stem from the blurred boundary between privacy-related and general information, making it difficult for models to distinguish between them effectively. Furthermore, the high accuracy indicates that models can infer sensitive attributes not explicitly present in the context, such as a speaker’s likely ethnicity, based on indirect cues like accent, highlighting the risk of implicit privacy inference.
@@
Introducing prompts containing privacy protection content (prompt engineering) can significantly enhance the model’s ability to prevent direct privacy leaks and reduce the refusal leakage rate. For example, the Gemini series achieves over an 80% increase in refusal leakage rates for sensitive information such as bank account numbers and home addresses when enhanced prompts are used. However, this approach has limited effectiveness in mitigating inference-based privacy leaks and may even lead to a decrease in refusal rates. For instance, after introducing privacy-enhanced prompts, SALMONN experiences a 21% increase in leakage rate in age inference tasks.
@@
The experimental results in Table˜20 also reveal differences in privacy awareness between audio and text models. Similar to audio models, the text-based GPT-4o series demonstrates strong security awareness. However, overall, text models tend to have lower refusal rates, indicating slightly reduced sensitivity to privacy information compared to audio models. Nevertheless, through the application of prompt engineering techniques, the privacy protection capabilities of text models can still be significantly improved, although the improvement is typically not as substantial as that seen in audio models. For example, Gemini-2.5 Flash achieves an improvement of less than 20% in protecting social security number under enhanced prompting.
@@ H.2 Dataset Construction Method
To ensure the reliability and effectiveness of the native audio quality, we introduce big_bench_audio dataset (Srivastava et al., 2022) as one of the native voice data samples. The subsequent audio production follows the following process:
@@ H.3.3 Results
| Qwen2.5-Omni | 8.14 | 7.10 | 7.50 | 7.93 | 7.12 | 7.17 |
| SALMONN | 2.00 | 6.42 | 4.57 | 2.94 | 7.16 | 6.66 |
| Ultravox | 4.00 | 7.53 | 7.30 | 6.53 | 6.70 | 7.00 |
| Step-Fun | 5.00 | 7.48 | 8.20 | 7.42 | 5.89 | 7.08 |
| Kimi Audio | 5.67 | 6.83 | 6.00 | 6.83 | 7.08 | 6.94 |
| Step-Audio2 | 6.18 | 6.58 | 7.92 | 6.82 | 0.00 | 6.78 |
| OpenS2S | 8.25 | 6.46 | 5.17 | 6.39 | 2.33 | 6.25 |
@@
| Gemini-1.5 Pro | 8.57 | 8.21 | 8.23 | 8.16 | 6.09 | 7.43 |
| Gemini-2.5 Flash | 8.16 | 8.38 | 8.28 | 7.93 | 6.36 | 7.76 |
| Gemini-2.5 Pro | 8.88 | 8.68 | 8.50 | 8.18 | 7.46 | 7.71 |
| GPT-4o Audio | 5.90 | 5.50 | 8.33 | 7.31 | 7.62 | 6.27 |
| GPT-4o mini Audio | 8.33 | 6.90 | 7.69 | 6.00 | 5.77 | 7.25 |
| Average | 6.67 | 6.94 | 7.08 | 6.89 | 6.27 | 6.95 |
: AR: Adversarial Robustness; AQV: Audio Quality Variation; BC: Background Conversation; ES: Environmental Sound; MS: Multiple Speakers; NI: Noise Interference.
| Open-source Models | ||||||||
| Test Type | MiniCPM-o 2.6 | Qwen2-Audio | SALMONN | Ultravox | Step-Fun | OpenS2S | Kimi Audio | Qwen2.5-Omni |
| Adversarial Robustness | 8.27 | 6.06 | 5.84 | 1.00 | 7.12 | 1.57 | 1.42 | 5.76 |
| Audio Quality Variation | 8.56 | 5.90 | 6.25 | 1.29 | 7.06 | 1.39 | 4.10 | 6.17 |
| Background Conversation | 8.35 | 6.40 | 6.58 | 1.06 | 7.06 | 1.42 | 4.08 | 6.29 |
| Environmental Sound | 8.19 | 6.43 | 7.06 | 1.27 | 7.28 | 1.86 | 4.50 | 6.30 |
| Multiple Speakers | 8.74 | 6.78 | 6.33 | 2.44 | 7.22 | 3.14 | 2.03 | 7.67 |
| Noise Interference | 4.27 | 3.83 | 4.22 | 1.34 | 6.52 | 1.42 | 3.46 | 3.56 |
| Average | 7.73 | 5.90 | 6.05 | 1.40 | 7.04 | 1.80 | 3.26 | 5.96 |
| Closed-source Models | |||||
| Test Type | Gemini-1.5 Pro | Gemini-2.5 Flash | Gemini-2.5 Pro | GPT-4o Audio | GPT-4o mini Audio |
| Adversarial Robustness | 8.09 | 7.61 | 8.17 | 6.70 | 1.44 |
| Audio Quality Variation | 7.90 | 7.59 | 8.17 | 5.80 | 1.73 |
| Background Conversation | 7.71 | 6.87 | 7.35 | 6.93 | 1.73 |
| Environmental Sound | 8.06 | 7.03 | 7.50 | 6.72 | 2.36 |
| Multiple Speakers | 7.66 | 7.24 | 7.99 | 8.39 | 4.74 |
| Noise Interference | 5.86 | 5.61 | 6.37 | 2.85 | 1.67 |
| Average | 7.55 | 6.99 | 7.59 | 6.23 | 2.28 |
@@
We evaluate the robustness of nine models against various auditory challenges in Appendix H.3.1, with detailed results presented in Table 21 Table 22 Table 23 and Talbe 24. The results reveal the following key findings:
@@ I.2 Dataset Construction Method
Voice Cloning Spoofing (200 samples). We collected 100 samples of voice wake-up words from an existing dataset (He et al., 2022), with each sample consisting of approximately 2 seconds of wake-up word audio. These samples constitute the Voice Cloning Spoofing test set. Detailed wake-up words and their corresponding developers are listed in Table˜25. Each sample was cloned using F5-TTS (Chen et al., 2024), and we constructed paired audio clips composed of the original audio (used as the legitimate user’s enrollment input) and the cloned audio (used as the attack input attempting to bypass authentication). In addition, we designed two distinct authentication scenarios for the experiments—Awake Assistant and Open The Door. In the former, the wake-up words are used to authenticate an AI voice assistant, while in the latter, they are used to activate access to a high-security laboratory door, with an emphasis on the consequences of authentication failures. Through this experimental setup, we aim to explore how different system prompt formulations affect an ALLM’s ability to defend against voice cloning spoofing under identical audio inputs.
@@
Note: All data in this table are adapted from (He et al., 2022).
@@ I.3 Additional Implementation Details
To ensure our evaluation closely mirrors real-world scenarios, distinct inference prompts were designed for each data point across all experiments, complementing the audio inputs to the ALLMs. For the hybrid spoofing dataset, in particular, prompts were varied based on the specific attack scenario. Illustrative examples of these inference prompts are provided below.